OpenStudy (anonymous):

Select the first five terms in the arithmetic sequence an = 4n, starting with n = 1.

OpenStudy (neer2890):

an=4n so 1st term is a1=4*1=4 2nd term is a2= 4*2=8 .... ... 5th term is a5=4*5=20

OpenStudy (goformit100):

Welcome to Open Study Ma'am @fairyangel33

OpenStudy (neer2890):

you didn't welcome me.. @goformit100 .... lol

OpenStudy (goformit100):

Welcome back Sir @neer2890 ;) to OpenStudy

OpenStudy (neer2890):

haha... just joking dear... @goformit100

OpenStudy (goformit100):

Thank god Now I don't have any fear.

OpenStudy (neer2890):

in today's world, nobody fears from anybody....! how could you :)

OpenStudy (goformit100):

I am from India. I am fear of the new Railway ticket cost.

OpenStudy (neer2890):

ache dino ka wada hua tha brother,, lagta h aane wale hain..lol